# Distributions of Ulam Words up to Length 30

### v1

@@ 3. Unconditional Results
It is easy to check that the only elements in both sets with $y=1$ are $(1,1)$ and $(2,1)$. Our goal is to show that the iterative process for producing points in $\mathcal{S}$ with larger $y$ values is the same as for the first set. To do so, we will use a modified version of Pascal’s triangle, achieved by aligning all entries to the left and then rotating the resulting image counterclockwise 90 degrees.
@@
This is not the first time that Pascal’s triangle and the Sierpiński triangle has showed up in the study of Ulam words. As mentioned previously, Mandelshtam found an analogous pattern for words $0^{x}10^{y}1^{z}$ [9]; even before that, Bade et al. proved that the number of words of length $n$ with one $1$ is the $n$-th term in Gould’s sequence[1]—that is, the number of elements in the $n$-th row of Pascal’s triangle.
@@
Fix $x\in\mathbb{Z}_{\geq 2}$; let $L_{x}$ be the set of $y\in\mathbb{Z}_{\geq 1}$ such that $0^{y}1^{x-y}\in\mathscr{U}$. Then #$\#L_{x}$ is the $(x-1)$-th term of Gould’s sequence.
@@
We close by giving one more result, which is analogous to the following theorem due to Bade et al.
Theorem 3.7 (Theorem 3.6 in [1]).
For any $a,b\in\mathbb{Z}$ such that $a+b\geq 2$, $0^{a}1010^{b}\in\mathscr{U}$ if and only if $a+b=1\mod 2$.
@@ 4. Density and the Distribution of the Gaps
In their paper, Bade et al. conjectured that the density $\rho(n)$ of Ulam words of length $n$—as defined in Section 1—converges to some real number $10<r<$ 1 as $n\rightarrow\infty$ [1]. With our enlarged data set, we conjecture that $\rho(n)\rightarrow 0$; specifically, $\rho(n)=\Theta\left(n^{-3/10}\right)$. We originally conjectured this due to illustrations such as Figure 1, but there is an even more striking way to see this, in terms of the average gap.
Let $u_{1}<u_{2}<\ldots<u_{k_{n}}$ be the (ordered) elements of $\pi(\mathscr{U}_{n})$—that is, the integers that are images of the Ulam words of length $n$. The gaps between them are the consecutive differences $g_{1}=u_{2}-u_{1}$, $g_{2}=u_{3}-u_{2}$, and so on. It is easy to see that $u_{1}=\pi(0000\ldots 01)=1$ and $u_{k_{n}}=\pi(1111\ldots 10)=2^{n}-2$. Ergo, the average gap between Ulam words of length $n$ is
@@
from which we conclude that $\rho(n)^{-1}\asymp\mu_{g}(n)$. So, we must conclude that if $\rho(n)$ converges to a non-zero constant, then $\mu_{g}(n)$ must be bounded. But this does not appear to be the case, as evidenced by Figures 7 and 8. Instead, we conjecture the following.
@@
Conjecture 4.1.
As $n\rightarrow\infty$, $\mu_{g}(n)=\Theta(n^{3/10})$—indeed, it may be that there is a constant $c\approx 1.9$ such that $\mu_{g}(n)=cn^{3/10}+o(1)$. Consequently, $\rho(n)=\Theta(n^{-3/10})$.
Figure 9. A plot of the maximal gap between words of length $n$ for $2\leq n\leq 30$, and $f(n)=1.35^{n}$.
It is obvious that any gap is at least $1$ (in fact, it is an easy exercise to show that a gap of $1$ is always attained); on the other hand, numerical evidence suggests that the maximal gap grows exponentially—specifically, it is $O(r^{n})$ for some constant $r\approx 1.35$ (see Figure 9). We can also study the distribution of the gaps more generally, and a very curious phenomenon emerges: it appears that these distributions converge to some limiting distribution as $n\rightarrow\infty$, as described in Conjecture 1.3. A large number of these distributions are illustrated in Figure 1.3.
@@
There are some curious details to this apparent distribution. The first is that there is a clear bias against gaps that are either $2$ or $4$ modulo $6$. This is very odd, considering that we conjecture that Ulam words are equidistributed modulo $6$ in the limit. (See Section 5 for more about this.) The second is that just as the average gap appears to grow without bound, so does the standard deviation. However, the standard deviations are quite small—here is a table of the last few we were able to compute:
$\begin{array}[]{l|l}n&\text{Standard Deviation}\\ \hline\cr 21&6.090434611011227\\ 22&6.5739141205506035\\ 23&6.951985360977966\\ 24&7.48652894820725\\ 25&7.954513792250833\\ 26&8.410261056967098\\ 27&8.838421071953993\\ 28&9.345660478229291\\ 29&9.940553024054223\\ 30&10.500749743720565\end{array}$
@@ 5. Modular Distribution
While there must always exist for any $n$ two congruence classes $a,b\in\mathbb{Z}/3\mathbb{Z}$ such that $\rho_{a,3}(n)=\rho_{b,3}(n)$, there is no reason why the last congruence class $c$ should be roughly equal. Indeed, for $n\leq 5$, $\rho_{c,3}(n)=0$. However, as Figure 11 shows, for larger $n$, it does appear to be the case that $\rho_{c,3}(n)\rightarrow\rho_{a,3}(n)=\rho_{b,3}(n)$.
Figure 11. A plot of the relative densities modulo $6$ as functions of the word length.
To help measure the extent to which words are equidistributing modulo $N$, we define the modular discrepancy.
@@
Trivially, saying that Ulam words equidistribute modulo $N$ is equivalent to saying that $d_{N}(n)\rightarrow 0$ as $n\rightarrow\infty$. Moreover, by appealing to the Chinese remainder theorem, proving that $d_{N}(n)\rightarrow 0$ for all $N$ is reducible to prove that $d_{p^{k}}(n)\rightarrow 0$ for all prime powers $p^{k}$. To investigate Conjecture 1.1, we computed $d_{p^{k}}(n)$ for all prime powers $p^{k}<30$—as near as we can tell, $d_{p^{k}}(n)$ decays exponentially as a function of $n$ (see Figure 12.)
@@
6. Description of the Algorithm
Our code for generating Ulam words and a list of all Ulam words up to length $20$ are available on GitHub [^1] https://github.com/asheydva/Ulam-Words.git. Below, we describe the algorithm, the current roadblocks, and possible theoretical improvements for the benefit of the reader.
One of the key components to the fast performance of the algorithm is the implementation of hash functions. Specifically, we made use of Python’s ‘set’ data type, which uses hash tables to make searching through the set for an element take $\Theta(1)$ operations on average. We use this as follows: all Ulam words that have been computed during previous iterations are stored in a dictionary UlamSet such that for any integer $k$, $\text{UlamSet}(k)$ returns the images of Ulam words of length $k$ under the map $\pi$, stored as a set. Ulam words are stored in terms of their associated integers to reduce memory requirements.
Aside from the dictionary UlamSet, we will also need a command we shall call Split—given a word $w$ and an integer $k$, it splits $w$ into two subwords by taking the first $k$ letters of $w$ as one word and the remaining letters as the second word. The specifics of how to implement this function are, of course, language-dependent, but it should never take more than $O(n)$ operations, where $n$ is the length of $w$. With this, we can define a function to test whether words of length $n$ are Ulam or not, assuming we have already determined the Ulam words of length $11\leq k\leq n-$ 1.
Algorithm 6.1.
On an input of a word $w$ of length $n$, outputs whether it is Ulam or not.
1:procedure UlamWordQ($w$)
[…]
3:     for $11\leq k\leq n-$ 1 do
4:         $w_{1},w_{2}\leftarrow\text{Split}(w,k)$
5:         $m_{1}\leftarrow\pi(w_{1})$
6:         $m_{2}\leftarrow\pi(w_{2})$
7:         if $m_{1}\in\text{UlamSet}(k)$ and $m_{2}\in\text{UlamSet}(n-k)$ and $m_{1}\neq m_{2}$ then
[…]
10:              break
[…]
Remark 6.1.
Here, the command ‘break’ functions as it does in Python: if it is ever activated, the algorithm exits the for-loop that it is currently executing.
Proof of Correctness.
[…]
Theoretically, the worst-case running time for UlamWordsQ is $O(n^{2}\#\mathscr{U}_{n-1})$ and thus, conjecturally, $O(n^{17/10}2^{n})$. However, in practice, the actual worst case seems to be $O(n^{2})$, which occurs whenever $w$ is Ulam or when it has no representations at all; the latter case appears to be rare. This is sufficiently good for our purposes.
With this, we just need one extra function: $\text{Binarize}(k,n)$, which returns the binary representation of $k$, padded with zeros on the left such that it has length $n$. Observe that listing all integers between $1$ and $12^{n}-$ 1 is then equivalent to listing all words of length $n$ (other than $0^{n}$ and $1^{n}$, which are never Ulam words).
Algorithm 6.2.
On an input of an integer $n$, prints all Ulam words of length $n$.
1:procedure UlamWords($n$)
2:     for $21\leq k\leq 2^{n}-$ 2 do
3:         $w\leftarrow\text{Binarize}(k)$
4:         if $\text{UlamWordQ}(w)$ then
5:              print $w$
Remark 6.2.
In practice, rather than printing, we instead write to a file. Moreover, we do not store $w$ but $k=\pi(w)$ to reduce memory requirements.
Proof of Correctness.
This is immediate. ∎
The running time of this algorithm should be something like $O(n^{2}2^{n})$, which is consistent with our observations. Since there are, conjecturally, $\approx n^{-3/10}2^{n}$ Ulam words of length $n$, there does not appear to be much room for improvement. The larger issue is memory. Because the algorithm requires storing the dictionary UlamSet in RAM, there is a hardware-dependent limit on the length of Ulam words that can be computed. On the personal computer that we used, we hit this limit at $n=30$—for context, the produced text file of all Ulam words of length $\leq 30$ was $4.11$ gigabytes.
The problem is imminently parallelizable: it is possible to test each word of length $n$ for Ulam-ness separately. Running it on a supercomputer with significantly more RAM would allow pushing past the limit that we encountered. Unfortunately, because of the exponentially growing time and memory requirements, even pushing to $n=40$ would pose substantial difficulties [^2] To give some perspective, we would expect that simply storing all Ulam words up to that size would require over $16$ terabytes..
Acknowledgment.
Our collaboration was funded by four Bates College grants, all awarded by the Dean of Faculty’s office: a STEM Faculty-Student Summer Research Grant and three Summer Research Fellowships.

### v2

@@ 4. Patterns in Ulam Words
It is easy to check that the only elements in both sets with $x\leq 2$ are $(1,1)$ and $(2,1)$. Our goal is to show that the iterative process for producing points in $\mathcal{S}$ with larger $x$ values is the same as for the first set. To do so, we will use a modified version of Pascal’s triangle, achieved by aligning all entries to the left and then rotating the resulting image counterclockwise 90 degrees.
@@
This is not the first time that Pascal’s triangle and the Sierpiński triangle have shown up in the study of Ulam words. As mentioned previously, Mandelshtam found an analogous pattern for words $0^{x}10^{y}1^{z}$ [9]; even before that, Bade et al. proved that the number of words of length $n$ with one $1$ is the $n$-th term in Gould’s sequence[1]—that is, the number of elements in the $(n-1)$-th row of Pascal’s triangle.
@@
Fix $n\in\mathbb{Z}_{\geq 2}$; let $L_{n}$ be the set of $a\in\mathbb{Z}_{\geq 1}$ such that $0^{a}1^{n-a}\in\mathscr{U}$. Then #$\#L_{n}=G(n)$.
@@
We close by giving one more result, which is analogous to Theorem 3.3.
@@ 5. Density and the Distribution of the Gaps
We begin by proving that there is a relationship between the size of the average gap between consecutive Ulam words of length $n$, and the density of Ulam words.
Proof of Theorem 1.4.
Let $u_{1}<u_{2}<\ldots<u_{k_{n}}$ be the (ordered) elements of $\pi(\mathscr{U}_{n})$—that is, the integers that are images of the Ulam words of length $n$. The gaps between them are the consecutive differences $g_{1}=u_{2}-u_{1}$, $g_{2}=u_{3}-u_{2}$, and so on. It is easy to see that $u_{1}=\pi(0000\ldots 01)=1$ and $u_{k_{n}}=\pi(1111\ldots 10)=2^{n}-2$—this follows from Theorem 3.1, for example. Ergo, the average gap between Ulam words of length $n$ is
@@
from which we conclude that $\rho(n)^{-1}\asymp\mu_{g}(n)$. (We already know from Theorem 1.1 that $k_{n}\rightarrow\infty$ as $n\rightarrow\infty$.) In particular, if Conjecture 1.2 is true and $\mu_{g}(n)\asymp n^{3/10}$, then $\rho(n)\asymp n^{-3/10}$, which is Conjecture 1.1. ∎
Notice in particular that for $\rho(n)$ to converge to a non-zero constant, it must be that $\mu_{g}(n)$ is bounded. But this does not appear to be the case, as evidenced by Figures 6 and 7. Instead, as near as we can tell, the order of growth seems to be around $n^{3/10}$. With this in mind, it is perhaps worthwhile to examine the distribution of the gaps more closely.
@@
Figure 8. A logarithmic plot of the maximal gap between words of length $n$ for $2\leq n\leq 30$, and $f(n)=1.35^{n}$.
It is obvious that any gap is at least $1$ (in fact, it is an easy exercise to show that a gap of $1$ is always attained); on the other hand, numerical evidence suggests that the maximal gap grows exponentially—specifically, it is $O(r^{n})$ for some constant $r\approx 1.35$ (see Figure 8). This is a little strange, in that one would not immediately guess this looking at the probability distributions, as in Figure 9.
@@
Indeed, there are some curious details to this apparent distribution. The first is that there is a clear bias against gaps that are either $2$ or $4$ modulo $6$. This is very odd, considering that we conjecture that Ulam words are equidistributed modulo $6$ in the limit. (See Section 6 for more about this.) The second is that just as the average gap appears to grow without bound, so does the standard deviation. However, the standard deviations are quite small, of order around $n/3$—here is a table of the last few we were able to compute:
$\begin{array}[]{l|l}n&\text{Standard Deviation}\\ \hline\cr 21&6.09043461\\ 22&6.57391412\\ 23&6.95198536\\ 24&7.48652894\\ 25&7.95451379\\ 26&8.41026105\\ 27&8.83842107\\ 28&9.34566047\\ 29&9.94055302\\ 30&10.5007497\end{array}$
@@ 6. Modular Distribution
While there must always exist for any $n$ two congruence classes $a,b\in\mathbb{Z}/3\mathbb{Z}$ such that $\rho_{a,3}(n)=\rho_{b,3}(n)$, there is no reason why the last congruence class $c$ should be roughly equal. Indeed, for $n\leq 5$, $\rho_{c,3}(n)=0$. However, for larger $n$, it does appear to be the case that $\rho_{c,3}(n)\rightarrow\rho_{a,3}(n)=\rho_{b,3}(n)$; as we will illustrate presently. To help measure the extent to which words are equidistributing modulo $N$, we define the modular discrepancy.
@@
Trivially, saying that Ulam words equidistribute modulo $N$ is equivalent to saying that $d_{N}(n)\rightarrow 0$ as $n\rightarrow\infty$. Moreover, by appealing to the Chinese remainder theorem, proving that $d_{N}(n)\rightarrow 0$ for all $N$ is reducible to proving that $d_{p^{k}}(n)\rightarrow 0$ for all prime powers $p^{k}$. To investigate Conjecture 1.3, we computed $d_{p^{k}}(n)$ for all prime powers $p^{k}<30$—as near as we can tell, $d_{p^{k}}(n)$ decays exponentially as a function of $n$ (see Figure 10.)
@@
Acknowledgments.
Our collaboration was funded by four Bates College grants, all awarded by the Dean of Faculty’s office: a STEM Faculty-Student Summer Research Grant and three Summer Research Fellowships. We would also like to thank Tomás Oliveira e Silva for confirming our computations of #$\#\mathscr{U}_{n}$ and giving many helpful suggestions for improving the exposition.